1)integrate the following
: ∫4x3xdx
2)a number is chosen randomly from numbers 1 to 60.the probability that the chosen number is a multiple of 2 or 5 is
 

Grade:10

2 Answers

Aditya Gupta
4 years ago
1. Note that 4^x.3^x= 12^x
So the integral is M= ∫12^xdx
Substitute y=12^x so that dy=12^x.ln12 dx
So M= ∫dy/ln12
= y/ln12
= 12^x/ln12
2. Number of numbers that are multiples of 2 or 5=
Number of multiples of 2 + number of multiples of 5 - number of multiples of 10 
= 30 + 12 - 6
= 36
Hence probability= 36/60
= 3/5
KINDLY APPROVE :))
Arun
4 years ago
Dear student
 
1. integration of 12^x dx
 = 12^x / (ln12) + const
 
Hope it helps
 
PS-please ask only one question in one thread.